OpenStudy (anonymous):

Please help me with my homework i really dont know how to do these its algerbra please explain how to do these because i have slow mind so please explain each problem thanks 1. Is it possible for two linear functions whose graphs are parallel lines to have the same y intercept? explain. 2. Estimate the slope of a line that is perpendicular to the line through (2.07,8.95) and (-1.9,25.07). 3. Explain in words how tp write an equation in slope-intercept form that describes a line parallel to y-3=-6(x-3).

OpenStudy (anonymous):



OpenStudy (anonymous):

yes plaese the other problems to @Ck[-.-]

OpenStudy (anonymous):

I don't know about 2. But for 3: First solve the equation into y=mx+b y-3=-6(x-3) Distribute: y-3=-6x+18 y=-6x+21 Parallel means that they have the same slope. So a line that would be parallel would be y=-6x+(whatever. but not 21)